OpenStudy (anonymous):

If cosr=3/5 and sint=5/13, r and t in Q1, find cos(r+t)

OpenStudy (michele_laino):

hint: \[\Large \cos \left( {r + t} \right) = \cos r\cos t - \sin r\sin t\]

OpenStudy (anonymous):

would it be 16/65?